ganeshie8 (ganeshie8):

For all positive integers \(n\), \show that the decimal expansion of \(\large \dfrac{1}{n}\) terminates if and only if \(\large n = 2^a5^b\) \(a,b\ge 0\)

OpenStudy (alexandervonhumboldt2):

hmm so n needs to be >1 right?

ganeshie8 (ganeshie8):

yes assume n is positive integer :)

OpenStudy (alexandervonhumboldt2):

anyways if a abd b>0 then 2^a*5^b>0 too

OpenStudy (kainui):

No, not necessarily, there's no reason why a and b can't both be 0.

OpenStudy (alexandervonhumboldt2):

oh yeah

OpenStudy (alexandervonhumboldt2):

hmm is a0 and b=0 then 2^a*5^b=1

OpenStudy (alexandervonhumboldt2):

if a=0 and b=0 then 2^a*5^b=1

OpenStudy (kainui):

Yeah, so as long as we are calling that 1.0 it is the most rapidly terminating decimal of all of them... (or maybe technically a counter example)

OpenStudy (alexandervonhumboldt2):

but le me think

OpenStudy (alexandervonhumboldt2):

there cant be a case where 1/0. thats impossible

OpenStudy (kainui):

Since a,b are greater than or equal to 0, this is the minimum. To get zero in the denominator we would need negative infinity to be in the domain for a and b. \[\Large \frac{1}{2^05^0}=1 \]

OpenStudy (alexandervonhumboldt2):

n cannot be 0

OpenStudy (alexandervonhumboldt2):

2^0*5^0=1 thus n>0

ganeshie8 (ganeshie8):

right, let me add that to the main q so it becomes more clear :)

OpenStudy (alexandervonhumboldt2):

then 1/n<=1

OpenStudy (alexandervonhumboldt2):

\[1/n \le1\]

OpenStudy (alexandervonhumboldt2):

@ganeshie8 thats hard

OpenStudy (alexandervonhumboldt2):

so we just got that 1/n<=1 but what this gived us? hmmmmm

OpenStudy (alexandervonhumboldt2):

\[0<1/n \le1\]

ganeshie8 (ganeshie8):

below doesnt terminate ``` 1/3 = 0.3333333... ``` below terminates ``` 1/2 = 0.5 ```

OpenStudy (alexandervonhumboldt2):

hmmmmm gsh

OpenStudy (alexandervonhumboldt2):

@dan815

OpenStudy (kainui):

1/2=.5 and 1/5=.2 and are terminating decimals, and I can't see a reason why you would believe multiplying two nonrepeating decimals would create an infinitely repeating decimal. It seems to work because 2 and 5 are the only divisors of 10. I guess this probably doesn't count as a real proof.

ganeshie8 (ganeshie8):

Ahh nice that is a legitimate proof in one direction xD what about the converse ?

OpenStudy (kainui):

Well, 2 and 5 are the only divisors of 10 so any other choice won't work haha, I guess that's also sort of weak.

ganeshie8 (ganeshie8):

\(\Rightarrow\) 1/2 = .5 terminates so (1/2)^a terminates 1/5 = .2 terminates so (1/5)^b terminates so their product 1/(2^a5^b) also terminates

ganeshie8 (ganeshie8):

yeah for the converse, if there is some easy intuition of why only the combinations of divosors of 10 will work..

OpenStudy (kainui):

I guess my alternate thing was playing with this, if b>a then \[\Large \frac{1}{2^a5^b}=\frac{2^{b-a}}{10^b}\] and it should terminate since it divides evenly haha I don't know

OpenStudy (kainui):

Haha I don't really know how to prove things, I mean if it's true, it's true and there are many ways you could not believe that this is true haha

ganeshie8 (ganeshie8):

that looks more neat \[\Large \frac{1}{2^a5^b}=\frac{2^{b}5^a }{10^{a+b}} \] still this wont help in proving converse i think...

ganeshie8 (ganeshie8):

converse : \(\dfrac{1}{n}\) terminates \(\implies \) \(n = 2^a5^b\) this has to address other primes also i guess

OpenStudy (kainui):

Well powers of 2 and 5 are the only numbers that are evenly divided by powers of 10, everything else is relatively prime so nothing else will work. I guess that's really all I am saying is my proof of the converse but I guess that's not quite enough hmmm

ganeshie8 (ganeshie8):

you're saying \(n \nmid 10^a\) when \(n\) is made up of other primes ?

ganeshie8 (ganeshie8):

that looks good enough to me :)

ganeshie8 (ganeshie8):

i believe all other proofs boil down to saying the same thing. .

OpenStudy (kainui):

Haha so is that it? I feel like I kinda avoided stuff, like why am I not expected to prove that two nonrepeating decimals always produces a nonrepeating decimal haha. I don't know how to prove such a thing, but I wouldn't see a reason to. It would be like proving two integers can never multiply together to make an irrational number or something haha.

ganeshie8 (ganeshie8):

i think we take them for granted in number theory, we wry about such petty stuff only in analysis :P

OpenStudy (kainui):

Ok good, but yeah even so, I feel like there's sort of an endless pit of possible things we could prove, so I don't know what's exactly acceptable in the world of number theory XD